OpenStudy (anonymous):

Use the appropriate rule or combination of rules to find the derivative of the function defined below. y = sin (x^4) cos (x^4)

OpenStudy (anonymous):

i used the product rule for both sin and cos to get (sin*4x^3+x^4*cos)(cos*4x^3*x^4-sinx)

OpenStudy (anonymous):

but my answer should look something like: 4cos(x^4)^2*x^4 -4sin(x^4)^2*x^4

OpenStudy (anonymous):

product rule and chain rule

OpenStudy (anonymous):

first differentiate sin(x^4) and then x^4.. do the same for cos also

OpenStudy (anonymous):

the answer you wrote second is not right either you need \((fg)'=f'g+g'f\) with \[f(x)=\sin(x^4), f'(x)=\cos(x^4)\times 4x^3=4x^3\cos(x^4)\]

OpenStudy (anonymous):

similarly if \(g(x)=\cos(x^4)\) then \(g'(x)=-\sin(x^4)\times 4x^3=-4x^3\sin(x^4)\)

OpenStudy (anonymous):

@satellite73 said... there will be no x^4 term it is x^3 term

OpenStudy (anonymous):

i thought that sin and x^4 were two different factors.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

should i end up with something that looks like: \[4x^3*\cos(x^4)*\cos(x^4)+(-4sinx^3*\sin(x^4))*\sin(x^4)\]

OpenStudy (anonymous):

ok for simplicity.. do one thing.. let some y = x^4

OpenStudy (anonymous):

now your equation will be sin(y)cos(y)..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i get: siny'cosy+cosy'-siny

OpenStudy (anonymous):

you know how to differentiate this right.. that will (cos(y)^2-sin(y)^2)dy

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i do the product as: fg'+gf'

OpenStudy (anonymous):

yes perfect.. d(uv) =vd(u)+ud(v) right.. v =cos(y) u=sin(y).. from this you will get the above thing

OpenStudy (anonymous):

i don't see how you're getting it to the power of 2 when it should be 1-1. shouldnt it?

OpenStudy (anonymous):

now our y = x^4.. so y'= (4x^3) x'

OpenStudy (anonymous):

put y and y' in above equation.. what do you get.. 4x^3(cos(x^4)^2-sin(x^4)^2)

OpenStudy (anonymous):

i didn't get everything to the power of 2, i just had(sin(4x^3)+cos(x^4))+((cos(4x^3)-sin(x^4)

OpenStudy (anonymous):

what you have seems to look like that answer though is it?

OpenStudy (anonymous):

ok.. how did you get 4terms?

OpenStudy (anonymous):

mistake you are would probably be this.. you are combining both product rule and chain rule..

OpenStudy (anonymous):

i think that is what i'm doing, i thought that is what i neeeded to do. so my answer should be: 4x^3(cos(x^4)^2-sin(x^4)^2)

OpenStudy (anonymous):

see when you differentiate sin(x^4).. first you sin derivative (cos(x^4)*(derivative of x^4))this whole thing is multiplied by cos(x^4).. do the same for cos(x^4) also

OpenStudy (anonymous):

yes i have that done.

OpenStudy (anonymous):

is 4x^3(cos(x^4)^2-sin(x^4)^2) what i should have? i find it easier to have a correct answer so i can look through my workings and see what i've done wrong.

OpenStudy (anonymous):

yes.. that is what you should get..

OpenStudy (anonymous):

can i write it as: 4x^3*cos(x^4)^2-sin(x^4)^2